OpenStudy (anonymous):

Part 1: Find the polynomial f(x) that has the roots of –2, 3 of multiplicity 2. (4 points) Part 2: Explain how you would verify the zeros of f(x). (4 points)

OpenStudy (abb0t):

\(f(x) = (x + 2)(x - 3)^2\)

OpenStudy (anonymous):

can u explain?

OpenStudy (abb0t):

i JUST did.

OpenStudy (anonymous):

To verify the zeros just solve this \[0 = (x+2)\] and \[0 = (x-3)^{2}\]